The sum of two numbers is 6 less than twice the first number.
Their difference is 10 less than four times the second number.

Answer by awang1996(50)

The sum of two numbers is 6 less than twice the first number.
Their difference is 10 less than four times the second number.

first number = x
second number = y

x+y=2x-6 this equals
y=x-6

x-y=4y-10 this equals
x=5y-10

substitute the second equation into the first equation
x=5y-10 --> y=x-6
y=5y-10-6 (simplify)
y=5y-16 (subtract 5y)
-4y=-16 (divide by -4)
y=4

substitute back into second equation
x=5y-10
x=5(4)-10
x=20-10
x=10

x=10 y=4